OpenStudy (anonymous):

x^2 + 12x+36=16 give exact and approx amount to 3 decimal places. How do you do this problem?

OpenStudy (anonymous):

The first thing you do when you are solving and there is an x^2 term involved is to make it equal to 0. x^2 + 12x + 20 = 0

OpenStudy (anonymous):

then i am supposed to factor it and set the answers = to zero and solve?

OpenStudy (anonymous):

In this case you are looking for factors of 20 that add to get 12 1 * 20 2 * 10 4 * 5 the pair that adds to get 12 is the pair you use.... 2 and 10 in this case. (x + 2)(x + 10) = 0 then you set each one equal to 0 and solve. x + 2 = 0 or x + 10 = 0 x = -2 or x = -10